OpenStudy (anonymous):

A card is selected at random from a well-shuffled ordinary deck consisting of 13 spades and 13 clubs (printed with black ink) and 13 hearts and 13 diamonds (printed in red ink). Three spades, three clubs, three hearts, and three diamonds are "face cards" - The probability that the card is a heart or is a face card is? a) 0.23 b)0.42 c)0.19 d)0.48 e)0.25

OpenStudy (anonymous):

I feel like the answer is e but im not exactly sure

OpenStudy (anonymous):

Add all of your cards together and divide.

OpenStudy (anonymous):

So 64 cards total and I divide that by the number of hearts, which is 13 and the total number of face cards which is 12?

OpenStudy (cwrw238):

probability of a heart = 13/52 and probability of a face card = 12/52

OpenStudy (cwrw238):

how did you get 64 total cards?

OpenStudy (cwrw238):

I thought a face card is a king, queen or Jack

OpenStudy (anonymous):

Ohh wait i added the total amount of face down cards as well

OpenStudy (cwrw238):

wait 3 of the face cards are hearts so P(hearts) = 13/52 and P(FaceCard)=9/52

OpenStudy (cwrw238):

now you have to find P(Heart or Face Card)

OpenStudy (anonymous):

oh my gosh, probability is so confusing =(( So instead of 13/52+12/52, which equals .48 I would do 13/52+9/52=.42?

OpenStudy (cwrw238):

yes